OpenStudy (anonymous):

what are the solutions of the equation x^2-4=0. use a graph of the related function. @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

factor a^2-b^2=(a-b)(a+b)=(x-2)*(x+2) so the root srae 2, -2

OpenStudy (anonymous):

thanks @AlexandervonHumboldt2 you really helped me out thats all i have for now but i will proabably be back later.